OpenStudy (anonymous):

Estimate the circumference. Use 3.14 or 22/7 to approximate π. A. 18.75 cm B. 19.63 cm C. 37.5 cm D. 39.25 cm plz help will fan and medal

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

sorry it always cuts of the fraction

OpenStudy (anonymous):

whats the radius?

OpenStudy (anonymous):

12.5 sorry

OpenStudy (anonymous):

C=2pi*r C=12.5*pi C=39.27 cm Approximately 39.25 cm using 3.14 as pi

OpenStudy (anonymous):

thanx so much What are the diameter and radius of a circle with a circumference of 28π m? A. d = 28 m r = 14 m B. d = 14 m r = 28 m C. d = 14 m r = 7 m D. d = 7 m r = 14 m

OpenStudy (anonymous):

circumference is 28pi?

OpenStudy (anonymous):

yes would it be A?

OpenStudy (anonymous):

that is correct cause 28 is 2r and r=d/2=14

OpenStudy (anonymous):

thanx Which is the best approximation for the circumference of a circular fish pond that is 24 ft in diameter? A. 38 ft B. 72 ft C. 75 ft D. 150 ft

OpenStudy (anonymous):

would this 1 be c?

OpenStudy (anonymous):

C=2pi*r d=2r

OpenStudy (anonymous):

d=24 given, so multiply d*pi=75.40 which is close to 75 ft

OpenStudy (anonymous):

ok thanx What is the radius of of a circle with a circumference of 32π yd? A. 4 yd B. 8 yd C. 16 yd D. 32 yd last 1

OpenStudy (anonymous):

r=32/2 r=

OpenStudy (anonymous):

c.16 yds

OpenStudy (anonymous):

thanx I got 100% :)

OpenStudy (anonymous):

glad to help